A brick has dimensions 16cm×8cm×4cm.find the ratio of its heights of centres of gravity,when it have maximum stability and minimum stability

Grade:7

1 Answers

Kishorkumar
8 years ago
Maximum stability:
Dimensions :Length=16cm or 8 cm:Breadth=16 cm or 8 cm :Height=4 cm
Let us consider the three edge points (16,0,0),(0,8,0),(0,0,4)
Then the Centre of Gravity id the MidPoint of the three points =(16+0+0 /3   ,0+8+0 /3 ,0+0+4 /3)
                                                                                                     =(16/3, 8/3, 4/3)
 
Minimum Stability:
Dimensions::Length=8 cm :Breadth=4 cm :Height=16 cm
again we consider the three edge points (8,0,0),(0,4,0),(0,0,16)
Centre of Gravity =(8+0+0 /3 ,0+4+0 /3 ,0+0+16 /3 )
                             =(8/3 ,4/3 ,16/3)
Height of the Brick is the coordinate along the Z-axis
Therefore the ratio of Heights when Max and Min stability is
                                    =4/3    ÷ 16/3
                                    =4/16
                                    =1:4
Hence Solved .